News by Date
News by Category
Solar System
Stars
White Dwarfs
Supernovas
Neutron Stars
Black Holes
Milky Way Galaxy
Normal Galaxies
Quasars
Galaxy Clusters
Cosmology/Deep Field
Miscellaneous
Press Resources
Status Reports
Press Advisories
Image Releases
Release Guidelines
Image Use Policy
NASA TV
Biographies/Interviews
Web Shortcuts
Chandra Blog
RSS Feed
Chandra Mobile
Chronicle
Email Newsletter
News & Noteworthy
Image Use Policy
Questions & Answers
Glossary of Terms
Download Guide
Get Adobe Reader
Chandra @ NASA
Visit the Chandra pages at the NASA portal
Image Use
Image Use Policy & Request Form
Guidelines for utilizing images, applets, movies, and animations featured in this Web Site.
Getting Hard Copies of Images
Ways to obtain photos, slides, etc of Chandra images.
Operations CXO Status Report

Friday 7 December 2001 10.00am EST

During the last week the loads were interrupted on 1 Dec due to an execution of the Science Instrument safing SCS 107. The SCS was executed by the OBC as a result of the EPHIN hardware fault flag being set. Analysis of both real-time and dump data showed no problems with the hardware and a procedure was run on Dec 3 to re-initialize the OBC EPHIN process and resume science. The initial engineering assessment points to a glitch in the data path between the EPHIN, EIO and OBC (possibly due to a single event upset) that resulted in the OBC failing to detect an expected EPHIN counter increment. Installation of a glitch counter will be considered as a possible action from the event.

A real-time procedure was completed on Dec 6 to uplink the momentum unloading dead-zone patch to the B-side OBC. This follows the loading of the A-side patch on Nov 12 and completes the patch process.

The schedule for the coming week is shown below and includes a Target Of Opportunity accepted on Dec 3. The TOO, XRF01113,0 is a Gamma Ray Burster and will be observed in coordination with a number of other telescopes including HST, Keck and the VLA. This weeks schedule also includes a series of observations of the asteroid 1998 WZ24.

------------------------------------------
Radiation Belts                      Dec 9
XRF011130               ACIS-I
RCW 38                  ACIS-I       Dec 10
1998 WT24 (3 obs)       ACIS-S       Dec 11
Vela Pwn                ACIS-S
Radiation Belts
SNR 1987A               ACIS-S       Dec 12
0208-512                ACIS-S       Dec 13
UM425                   ACIS-S
Radiation Belts                      Dec 14 
IC4296                  ACIS-S       Dec 15
PSR 0656+14 (2 obs)     ACIS-S
Betelgeuse              ACIS-I
3C78                    ACIS-S
BRI 2212-1626           ACIS-S       Dec 16
Cygnus Loop NE1         ACIS-S
2MASSJ171442+2602       ACIS-I
AGX J170730+2353        ACIS-I

------------------------------------------

All spacecraft subsystems continued to operate nominally.

Return to Status Reports